Terapix Star Formation Region IC 1396, © 2001 CFHT
Minutes for the terameeting of Monday 25 November 2013
by HJMCC - Updated November 27th, 2013

Minutes for the Terameeting of Monday 25 November 2013 at 10:00

Present: Henry, Nawfel, Patrick

-  Absent: Fred, Olivier


-  Henry will copy NB data to the public FTP area
-  Fred will make a list of old machines and their characteristics (with a view to eventually giving them away.
-  Patrick will update us on the status of TERAPIX2
-  Nawfel will stack CS82 data and compare to 2MASS

Stripe82 (Nawfel)

-  Navel has done the comparison with the 2MASS
-  Nawfel has now swarped the images and verified that all is OK.
-  There will be 100 patches with image to stack.
-  Still to do: run quality FITS in the stacked images.

Pagani WIRCAM images

-  Images OFF have been processed to stack.
-  Still need to do second pass.

UltraVISTA phase 3 preparation (Henry)

-  Phase 3 processing ongoing.
-  Should be finished for the next meeting
-  Henry doing documentation for ESO and preparing release
-  Bo asked for narrow-band images: Henry will copy them to the FTP area

VIKING data (all)

-  VIKING data is on new cluster: not sure if we can use the new cluster to reduce this data however
-  Nawfel will be finished this week: can we process the central 3 deg^2 on the old cluster? (maybe on raid2 / fcix9?

Status TERAPIX (Patrick)

-  mix7,mix8,mix9,mix10 still to be stopped.
-  certain MIX machines 2? have flaky RAID cards - could put these machines as scratch RAID card ?
-  Fred not present at meeting.


-  Problem with installation of system was difficult; now resolved
-  Should be installed
-  mix33 - 200 GB one of these two machines is instable
-  mix32 - 400 GB
-  Laurent D. away: should be back this week. No news to report otherwise
-  Next meeting with service informatique?


-  4keur RAM
-  2keur spare disks
-  plus extra stuff which has
-  One large machine has been ordered (fcix11 140 Tb)
-  No new machine added to TERAPIX1 since 24 months


-  All machines have been turned off with the exception of cosmix5 sl2s
-  Status of Goldman disk?

Next meeting: Monday 16 December 10:00 in 281

Site Map  -   -  Contact
© Terapix 2003-2011